OpenStudy (anonymous):

Find an equation of the tangent line to y = [2x]/[x+1] at the point (1,1)

OpenStudy (anonymous):

Are you familiar with derivatives?

OpenStudy (anonymous):

@sonalig

OpenStudy (anonymous):

yes, I understand how to find the slope of the line but I am not sure how to find the y-intercept

OpenStudy (anonymous):

So you know what derivatives are and how to use them right? @sonalig

OpenStudy (anonymous):

Find dy/dx at the point (1,1) - that's you're slope. Then you use the point-slope formula: \[y-y_0=m(x-x_0),\text{ where}\\ (x_0,y_0)=\text{given point}\\ m=\text{slope}\]